# Emergence of generic first-passage time distributions for large Markovian networks

### v1

@@ I Introduction
Biochemical processes can be described by time-continuous and space-discrete master equations [13, 14]. Reaching a certain state in such a network is called a first passage and the corresponding time is called the first-passage time (FPT) (Fig. 1(a)). Like the underlying state variable, it is a stochastic variable with a certain distribution. The statistical properties of first-passage times play a central role in physics [15, 16, 17, 18, 19, 20, 21, 22], chemistry [23, 24, 25, 26], biology [27, 8, 28, 29, 30, 31, 32, 33], and economics [34, 35, 36], where they characterize noise-driven event timing, barrier-crossing dynamics, and rare transitions in complex stochastic systems [37, 38, 39, 7].
@@
In the context of master equation models for kinetic proofreading, it has been observed before, both through analytical and numerical means, that on the timescale of the mean completion time, the FPT density tends to converge to one of two simple forms in the limit of large network size: either to a delta distribution, indicating quasi-deterministic behavior, or to an exponential distribution, indicating quasi-memoryless behavior [40, 41] (Fig. 1(b)). As pointed out by the authors, this phenomenon helps to explain the success of coarse-grained Markov chain models for complex systems. It also raises the question how much microscopic information can be inferred from macroscopic observations. In fact, it is common practice to infer system properties from experimental first-passage time measurements [42, 43, 44, 45, 46, 47, 48]. If only the delta distribution or the exponential distribution, both of which are fully characterized by their means, emerged in the limit of large networks, such inference methods would become impossible in the limit of large systems. In the study of complex networks, ignorance about the microscopic details of the network remains a substantial obstacle to understanding these systems [49, 50, 51]. Model independent results are therefore particularly desirable [52]. In this context, Godec and Metzler [53, 54] identified the separation of time scales between direct and indirect trajectories as a mechanism leading to universal FPT-asymptotics.
@@
Here we show how the two emerging limit distributions identified earlier for kinetic proofreading are in general related to the distribution of the eigenvalues of the generator matrix. We show that the key quantity is the sum of the inverse eigenvalues of the generator matrix. If infinitely many eigenvalues contribute comparably, a delta function emerges, while the exponential function results, when one value $\lambda_{1}^{-1}$ dominates (Fig. 1(c)). However, these two cases are not simple mirrors of each other, because the delta distribution requires stricter conditions than the exponential distribution. The deep connection between eigenspectra and properties of the FPT distribution - especially in the case where detailed balance holds for the steady state distribution, ensuring that the spectrum is real - has proven fruitful in the context of extreme value and time-average statistics [59, 60]. Here we analyze this connection using the interpretation of first-passage times of Markovian networks in terms of graph theory, which recently gained more interest in the context of first passage time problems [61, 62, 63, 64, 65, 66, 67, 68, 69]. In particular, we will use the fact that first passage time moments can be obtained from a decomposition of the Markovian network into spanning trees and two-tree spanning forests (Fig. 1(d)). We illustrate and validate our theoretical analysis by stochastic computer simulations using the Gillespie algorithm.
@@ II.3 First-passage time and graph theory
$\mathcal{L}{\mathbf{f}}(s)=(K^{T}-sI)^{-1}\mathbf{f}_{0}=\sum_{n=0}^{\infty}s^{n}(K^{T-1})^{n}\mathbf{1},$ (8)
@@ IV Discussion
It has been reported before for kinetic proofreading networks that two simple limiting distributions arise in the limit of large networks [40, 41]. This observation is surprising in the light of the immense structural complexity of biochemical reaction networks [93]. It calls for a more thorough mathematical explanation, since FPTs in Markovian networks correspond to so-called phase-type distributions, which are dense in the class of distributions with non-negative support [80]. Consequently, they can, a priori, approximate any possible shape of an FPT distribution arbitrarily well.
@@
The physical expectation that a forward bias generically leads to the deterministic limit [40, 41] is mathematically subtle, as shown in Sec. III.5. Instead, one can consider the mean residual life of the process: it long-time limit corresponds to the principal eigenvalue. So if the mean residual life vanishes for large times compared to the mean first-passage time, one can conclude that the eigenvalue ratio must also vanish. The result yields a formal condition for the deterministic limit in terms of a forward bias. While conceptually clear, this formulation is not directly tied to structural network properties, for instance the statistics of the transition rates.
Simulations of first-passage times in random networks with random rates confirm that the deterministic (delta-like) limit emerges robustly in networks with a forward bias. In these cases, the ratio of the principal eigenvalue contribution to the total spectral weight, vanishes with increasing system size, in agreement with the spectral analysis. Conversely, in reversible networks a backward bias generically leads to dominance of a single eigenvalue and hence to exponential first-passage statistics, again as predicted by the spectral analysis.
@@
Overall, we conclude that the universality of these limiting first-passage time distributions holds under clearly identifiable structural conditions. First, the large-network limit can yield universal behavior only if the process starts sufficiently far away from the target state. Otherwise, local non-macroscopic structures dominate. Thus, a condition such as the macroscopic forest condition is needed for universality. Second, irreversible transitions can break this universality: while a network composed of strongly connected microscopic subnetworks yields several delta peaks, if multiple macroscopic strongly connected components exist with irreversible connections between them, the resulting FPT distribution becomes non-generic. One of the simplest example would be a system with two macroscopic components, each individually leading to exponential FPTs, but connected by an irreversible transition. The total FPT distribution in the limit $N\to\infty$ then becomes a convolution of two exponentials — neither purely exponential nor deterministic. Finally, universality cannot be expected if arbitrarily large outgoing rates are allowed (compare the example in Supplement V).
Understanding how simple macroscopic laws emerge from the intricate dynamics of microscopic constituents lies at the heart of statistical physics. While this question was originally treated within the context of equilibrium systems, recent decades have seen remarkable progress in extending these principles to non-equilibrium settings [57, 99, 100, 58], such as living systems [101, 102]. The methods established here might be helpful also in the context of stochastic thermodynamics, because they are valid for arbitrary Markov networks.
Our results highlight that even in high-dimensional Markovian networks, the interplay between spectral properties of the generator and the underlying graph structure governs the emergence of simple limiting first-passage times. For large and complex networks, approaches based on spectral theory [59, 90] and complementary graph-theoretical formalisms [69, 67] therefore provide a powerful unified formalism for uncovering general mechanisms that so far remained hidden in the enormous microscopic parameter space. These insights set the stage for extending the present results to broader classes of stochastic systems and for exploring universal behavior beyond the regimes addressed here.
@@ Acknowledgements.
JBV thanks the German Academic Scholarship Foundation (Studienstiftung des Deutschen Volkes) for support. We also acknowledge support by the Max Planck School Matter to Life funded by the Dieter Schwarz Foundation and the Max Planck Society.

### v2

@@ I Introduction
Reaching a certain state in a stochastic system is called a first passage and the corresponding time is called the first-passage time (FPT) (Fig. 1(a)). Like the underlying state variable, it is a stochastic variable with a certain distribution. The statistical properties of first-passage times play a central role in physics [13, 14, 15, 16, 17, 18, 19, 20], chemistry [21, 22, 23, 24], biology [25, 8, 26, 27, 28, 29, 30, 31], and economics [32, 33, 34], where they characterize noise-driven event timing, barrier-crossing dynamics, and rare transitions in complex stochastic systems [35, 36, 37, 7].
@@
Previous work has established that FPT-statistics in confined and diffusion-controlled systems often exhibit robust and universal features. In particular, seminal work by Bénichou, Voituriez and collaborators [38, 15] showed that first-passage processes for diffusion in complex geometries can be classified in terms of transport properties, such as compact versus non-compact exploration, leading to universal scaling of mean FPTs and simple limiting distributions. Recent studies by Baravi, Kessler and Barkai [39, 40] showed that these FPT-distributions generically exhibit a characteristic biscaling structure with a short time dependence that reflects the initial condition and a universal long-time exponential tail. This is in agreement with the work of Godec and Metzler, who identified the separation of time scales between direct and indirect trajectories as a mechanism leading to universal FPT-asymptotics [41, 42].
Here we address the question of universal limits of FPT-distributions for time-continuous and space-discrete master equations [43, 44], which are often used to describe networks of biochemical processes. In the context of master equation models for kinetic proofreading, it has been observed before, both through analytical and numerical means, that on the timescale of the mean completion time, the FPT-density tends to converge to one of two simple forms in the limit of large network size: either to a delta distribution, indicating quasi-deterministic behavior, or to an exponential distribution, indicating quasi-memoryless behavior [45, 46] (Fig. 1(b)). As pointed out by the authors, this phenomenon helps to explain the success of coarse-grained Markov chain models for complex systems. It also raises the question how much microscopic information can be inferred from macroscopic observations. In fact, it is common practice to infer system properties from experimental first-passage time measurements [47, 48, 49, 50, 51, 52, 53]. If only the delta distribution or the exponential distribution, both of which are fully characterized by their means, emerged in the limit of large networks, such inference methods would become impossible in the limit of large systems. In the study of complex networks, ignorance about the microscopic details of the network remains a substantial obstacle to understanding these systems [54, 55, 56]. Model independent results are therefore particularly desirable [57].
@@
Here we show how the two emerging limit distributions identified earlier for kinetic proofreading are in general related to the distribution of the eigenvalues of the generator matrix. We show that the key quantity is the sum of the inverse eigenvalues of the generator matrix. If infinitely many eigenvalues contribute comparably, a delta function emerges, while the exponential function results, when one value $\lambda_{1}^{-1}$ dominates (Fig. 1(c)). However, these two cases are not simple mirrors of each other, because the delta distribution requires stricter conditions than the exponential distribution. The deep connection between eigenspectra and properties of the FPT distribution - especially in the case where detailed balance holds for the steady state distribution, ensuring that the spectrum is real - has proven fruitful in the context of extreme value and time-average statistics [62, 63]. Here we analyze this connection using the interpretation of first-passage times of Markovian networks in terms of graph theory, which recently gained more interest in the context of first-passage time problems [64, 65, 66, 67, 68, 69, 70, 71, 72]. In particular, we will use the fact that first-passage time moments can be obtained from a decomposition of the Markovian network into spanning trees and two-tree spanning forests (Fig. 1(d)). We illustrate and validate our theoretical analysis by stochastic computer simulations using the Gillespie algorithm.
@@ II.3 First-passage time and graph theory
$\mathcal{L}{\mathbf{f}}(s)=(sI-K^{T})^{-1}\mathbf{f}_{0}=\sum_{n=0}^{\infty}s^{n}(K^{T-1})^{n}\mathbf{1},$ (8)
@@ IV Discussion
It has been reported before for kinetic proofreading networks that two simple limiting distributions arise in the limit of large networks [45, 46]. This observation is surprising in light of the immense structural complexity of biochemical reaction networks [96]. It calls for a more thorough mathematical explanation, since FPTs in Markovian networks correspond to so-called phase-type distributions, which are dense in the class of distributions with non-negative support [83]. Consequently, they can, a priori, approximate any possible shape of an FPT distribution arbitrarily well.
@@
The physical expectation that a forward bias generically leads to the deterministic limit [45, 46] is mathematically subtle, as demonstrated in Sec. III.5. A more precise characterization can be obtained by considering the mean residual life of the process, whose long-time limit is governed by the principal eigenvalue. If the mean residual life becomes small at long times compared to the mean first-passage time, this implies that the contribution of the dominant eigenvalue vanishes, providing a formal condition for the deterministic limit. While this perspective is conceptually clear, it is not directly expressed in terms of structural properties of the network, such as the statistics of the transition rates.
This limitation becomes apparent already for one-step master equations, which can be interpreted as random walks in random environments. In this case, the relevant control parameter is not a simple average of the forward and backward rates, but rather the logarithmic bias $\langle\log(k_{i}/r_{i})\rangle$ [102]. Its sign determines the limiting behavior: for $\langle\log(k_{i}/r_{i})\rangle>0$, the dynamics approach the deterministic limit, whereas for $\langle\log(k_{i}/r_{i})\rangle<0$, they converge to the exponential limit. This illustrates that even in the simplest setting, naive bias measures are insufficient, and that identifying the correct effective quantity requires a more refined description.
Simulations of first-passage times in random networks with random rates confirm that the deterministic (delta-like) limit emerges robustly in networks with a forward bias. In these cases, the ratio of the principal eigenvalue contribution to the total spectral weight, vanishes with increasing system size, in agreement with the spectral analysis. Conversely, in reversible networks, a backward bias generically leads to dominance of a single eigenvalue and hence to exponential first-passage statistics, again as predicted by the spectral analysis.
@@
Overall, we conclude that the universality of these limiting first-passage time distributions holds under clearly identifiable structural conditions. First, the large-network limit can yield universal behavior only if the process starts sufficiently far away from the target state. Otherwise, local non-macroscopic structures dominate. Thus, a condition such as the macroscopic forest condition is needed for universality. Second, irreversible transitions can break this universality: while a network composed of strongly connected microscopic subnetworks yields several delta peaks, if multiple macroscopic strongly connected components exist with irreversible connections between them, the resulting FPT distribution becomes non-generic. One of the simplest examples would be a system with two macroscopic components, each individually leading to exponential FPTs, but connected by an irreversible transition. The total FPT distribution in the limit $N\to\infty$ then becomes a convolution of two exponentials — neither purely exponential nor deterministic. Finally, universality cannot be expected if arbitrarily large outgoing rates are allowed (compare the example in Supplement V).
Understanding how simple macroscopic laws emerge from the intricate dynamics of microscopic constituents lies at the heart of statistical physics. While this question was originally treated within the context of equilibrium systems, recent decades have seen remarkable progress in extending these principles to non-equilibrium settings [60, 103, 104, 61], such as living systems [105, 106]. Earlier work has made large progress to identify universal limiting distributions for space-continuous processes, often using Laplace-transforms [38, 15, 42, 39, 40]. Our work takes a complementary point of view by considering space-discrete Markovian networks and without assuming an underlying spatial or diffusive structure. In fact our results are valid also for different rates for each transition. We show that the emergence of simple FPT-distributions can be understood in terms of the eigenvalue distribution of the generator, providing a unifying classification that extends the concept of universality in first-passage processes from geometry- and transport-controlled kinetics to general networks with a large parameter space such as biochemical reaction networks. In this framework, exponential behavior arises from the dominance of a single eigenvalue, consistent with previous observations in diffusion-controlled systems. In contrast, when many eigenvalues contribute comparably, the FPT-distribution approaches a quasi-deterministic distribution around its mean, which we characterize here as a distinct generic limit.
More broadly, our results highlight that even in high-dimensional Markovian networks, the interplay between spectral properties of the generator and the underlying graph structure governs the emergence of simple limiting first-passage times. For large and complex networks, approaches based on spectral theory [62, 93] and complementary graph-theoretical formalisms [72, 70] therefore provide a powerful unified framework for uncovering general mechanisms that remain hidden in the enormous microscopic parameter space. These insights set the stage for extending the present results to broader classes of stochastic systems and for exploring universal behavior beyond the regimes addressed here.
@@ Acknowledgements.
JBV thanks the German Academic Scholarship Foundation (Studienstiftung des Deutschen Volkes) for support. We thank David Geldbach and Eli Barkai for helpful comments. We also acknowledge support by the Max Planck School Matter to Life funded by the Dieter Schwarz Foundation and the Max Planck Society.